# Tree Embedding in High Dimensions:

### v1

@@ 11    $\mu(v)\leftarrow$ arbitrarily match the points in $L(v)\setminus L^{\prime}(v)$
Any matching $\mu$ of points $P$ which satisfies lemma˜5.7 is of optimal cost with respect to tree metric $\mathcal{\tau}$.
@@
Fix an arbitrary non-root node $v\in T$ and its parent $u$. $\mu^{*}$ must match at least |A(v)B(v)||A(v)-B(v)| points of $P(v)$ with points not in $P$. Hence, $\Pi^{*}$ will contain edge $(u,v)$ at least |A(v)B(v)||A(v)-B(v)| times. Now consider $\Pi$. By the invariant it matches at most |A(v)B(v)||A(v)-B(v)| points of $P(v)$ with points not in $P(v)$. Hence, $\Pi$ contains edge $(u,v)$ exactly |A(v)B(v)||A(v)-B(v)| times. As edge weights are non-negative in $T$ this implies that $c(\mu)\leq c(\mu^{*})$ (and as $c(\mu^{*})$ is of minimum cost by assumption $c(\mu)=c(\mu^{*})$).
@@
In our implementation the output of this simple static algorithm can be maintained efficiently under the type of updates defined by Theorem˜4.2. First, observe that if we insert a path starting from a newly inserted leaf with a deactivated point to an existing node, in order to update the output of the algorithm, it is sufficient to set $L(v)$ and $L^{\prime}(v)$ values to be empty for each inserted node $v$.
We will describe a constant time operation how, starting from a leaf and walking to the root, we may handle point deactivations and point activations (Algorithm˜6 and Algorithm˜7). We say that node $v$ needs to handle a point insertion (or deletion) if for some child node $u$ of $v$ the set $L^{\prime}(u)$ undergoes a point insertion (or deletion). Both algorithms consist of procedures described in Algorithm˜8 and Algorithm˜9, in a bottom-up manner.
@@
First, assume, without loss of generality, that node $v\in T$ needs to handle a point insertion of $a\in A$ (see Algorithm˜8). First we need to extend $L(v)$ by $a$. If $A(v)\geq B(v)$ then before the insertion $L^{\prime}(v)$ was either empty or consisted points of $A$. In this case we may simply extend $L^{\prime}(v)$ by $a$ as well and ask the parent of $v$ to handle the insertion of $a$.
@@ 15 $L^{\prime}(v)\leftarrow L^{\prime}(v)\cup\{a\}$
We can describe an analogous process for the deletion of point $a$ from node $v$ (see Algorithm˜9). We first set $L(v)=L(v)\setminus\{a\}$. If $a\in L^{\prime}(v)$ then we can simply set $L^{\prime}(v)=L^{\prime}(v)\setminus\{a\}$ and ask the parent of $v$ to handle the deletion of $a$. Otherwise, if $A(v)>B(v)$ before the update we can select an arbitrary $a^{\prime}$ point of $L^{\prime}(v)$ and set $\mu(a^{\prime})=\mu(a)$, delete $a$ from $\mu$, set $L^{\prime}(v)=L^{\prime}(v)\setminus\{a^{\prime}\}$ and ask the parent of $v$ to handle the deletion of $a^{\prime}$. If $B(v)\geq A(v)$ before the deletion then let $b=\mu(a)$, we, delete the value of $a$ from $\mu$, set $L^{\prime}(v)=L^{\prime}(v)\cup b$ and ask the parent of $v$ to handle the insertion of $b$.
@@
Dynamic algorithms Algorithm˜6 and Algorithm˜7 maintain the solution to the Euclidean bipartite matching w.r.t. tree metric $\mathcal{T}$.
@@
We prove the above claim by showing that the invariant Lemma˜5.7 is maintained after every update.
@@ 14 $L^{\prime}(v)\leftarrow L^{\prime}(v)\setminus\{a\}$
Corollary˜5.5 now follows directly from Lemma˜5.9 and Lemma˜5.10 and Theorem˜4.2.
@@ 5.3 Geometric Transport
In this section we describe our implementation of the fully dynamic geometric transport problem. The algorithm is mostly identical to that of section˜5.2 for Euclidean bipartite matching. However, in the case of geometric transport we allow for point pairs with large weights to be inserted and deleted from the input, hence the recourse of the actual assignment describing the optimal solution can be large after each update. Therefore, our algorithm only maintains the cost of the optimal solution.
@@
Similarly to our other applications, the algorithm can either obtain $O_{\epsilon}(\log n)$-approximation in $\tilde{O}(n^{\epsilon}+d)$ update time or $\tilde{O}(1)$-approximation in $\tilde{O}(d)$ update time. Our dynamic implementation will maintain the exact cost of an optimal solution with respect to the tree embedding described by theorem˜4.2. Hence, corollary˜5.11 is a direct implication of lemma˜4.4 and lemma˜5.12.
@@
First, notice that we can trivially turn an instance of geometric transport into an instance of Euclidean bipartite matching through creating $w(v)$ copies for all leaves $v$ in $T$. Therefore, it is enough to show that the cost of the matching $\mu$ produced by the static algorithm Algorithm˜5 of section˜5.2 is exactly $\tilde{\mu}=\sum_{v\in T}||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$.
Now fix some arbitrary optimal matching $\mu^{*}$ on this Euclidean matching instance. Let $\Pi$ and $\Pi^{*}$ correspond to the multi-set of edges described by the union of paths on $T$ connecting the matched points of $\mu$ and $\mu^{*}$ respectively. Fix an arbitrary edge between $v$ and its parent $u$. In $\Pi^{*}$ this edge must appear at least $||A_{w}(v)|-|B_{w}(v)||$ as that many points of $P(v)$ cannot be matched to each other. Similarly, due to lemma˜5.7 this edge will appear at most $||A_{w}(v)|-|B_{w}(v)||$ times.
The paths in $\Pi$ crossing this edge contribute at most $||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$ to $c(\mu)$. Similarly, the paths in $\Pi^{*}$ crossing this edge due to crossing it contribute at least $||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$ to $c_{T}(\mu^{*})$ (as $\mu^{*}$ might not match $\min\{|A_{w}(v)|,|B_{w}(v)|\}$ pairs from $P(v)$). Summing over all $v\in T$, we get $c(\mu)\leq\tilde{\mu}\leq c(\mu^{*})$. As $c(\mu)\geq c(\mu^{*})$, we finally conclude that $c(\mu)=c(\mu^{*})=\sum_{v\in T}||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$.
@@ 5.4 Euclidean Matching
In this subsection we describe our dynamic algorithm for dynamic Euclidean matching. The algorithm is similar to that of section˜5.2 for the bipartite version of the problem.
@@
We will maintain a tree embedding of the input points using theorem˜4.2. Let $T$ be the $2$-HST maintained by the algorithm, where leaves correspond the points of $P$. We will maintain an exact solution to the problem with respect to the distance metric defined by $T$. Then corollary˜5.14 is a direct implication of lemma˜4.4 and lemma˜5.15.
@@ 6 $L^{\prime}(v)\leftarrow\emptyset$
Observe that the matching $\mu$ computed by the static algorithm satisfies ˜5.16 by construction. Now we will show that this condition is sufficient to prove optimality.
@@
If matching $\mu$ satisfies ˜5.16 then $\mu$ is a minimum cost matching of $P$ with respect to the tree metric $\mathcal{\tau}$.
@@
Consider an arbitrary node $v\in T$ with parent $u$. If $|P(v)|$ is odd then $\mu^{*}$ must match a point of $P(v)$ with some point outside of $P(v)$. Hence $\Pi^{*}$ contains $(u,v)$. By assumption $\mu$ only matches one point of $P(v)$ outside of $P(v)$, hence $\Pi$ contains $(u,v)$ once. If $|P(v)|$ is even then, $\mu$ matched all points of $P(v)$ with each other, hence $(u,v)\notin\Pi$. This implies that $c_{T}(\mu)\leq c_{T}(\mu^{*})$.
@@
We will now describe how to dynamically maintain a matching $\mu$ satisfying ˜5.16 under Type 1) and Type 2) updates in worst-case $\tilde{O}(1)$ update time. For sake of simplicity we will assume that $|P|$ is even at all times, that is updates come in pairs.
@@ 17    $\textbf{Delete}(u,p^{\prime\prime})$
The recursion follows a leaf to root path and takes $O(1)$ time to execute for a given note hence runs in $O(m)=\tilde{O}(1)$ time. Furthermore, it ensures that from $P(v)$ only the up to one point in $L^{\prime}(v)$ is not matched with points in $P(v)$. Hence, the algorithm satisfies ˜5.16 and by ˜5.17 it maintains a minimum cost matching of $P$ with respect to the tree metric $\mathcal{\tau}$.
Appendix A Local Queries for Hash Functions
In Lemma˜4.4, we state that the $(\Gamma,\Lambda)$-hash $\varphi$ with diameter bound $\tau$ can evaluate $\varphi(B(p,\tau/\Gamma))$ in $O(\operatorname{poly}(d)\cdot|\varphi(B(p,\tau/\Gamma))|$ time. However, this part of the statement did not appear in [FJL+25], and here we provide a proof sketch.

### v2

@@ 10    $\mu(v)\leftarrow$ arbitrarily match the points in $L(v)\setminus L^{\prime}(v)$
Any matching $\mu$ of points $P$ which satisfies Lemma 5.7 is of optimal cost with respect to tree metric $\mathcal{\tau}$.
@@
Fix an arbitrary non-root node $v\in T$ and its parent $u$. $\mu^{*}$ must match at least |A(v)B(v)||A(v)-B(v)| points of $P(v)$ with points not in $P$. Hence, $\Pi^{*}$ will contain edge $(u,v)$ at least |A(v)B(v)||A(v)-B(v)| times. Now consider $\Pi$. By the invariant it matches at most |A(v)B(v)||A(v)-B(v)| points of $P(v)$ with points not in $P(v)$. Hence, $\Pi$ contains edge $(u,v)$ exactly |A(v)B(v)||A(v)-B(v)| times. As edge weights are non-negative in $T$ this implies that $c(\mu)\leq c(\mu^{*})$ (and as $c(\mu^{*})$ is of minimum cost by assumption $c(\mu)=c(\mu^{*})$). ∎
@@
In our implementation the output of this simple static algorithm can be maintained efficiently under the type of updates defined by Theorem 4.2. First, observe that if we insert a path starting from a newly inserted leaf with a deactivated point to an existing node, in order to update the output of the algorithm, it is sufficient to set $L(v)$ and $L^{\prime}(v)$ values to be empty for each inserted node $v$.
We will describe a constant time operation how, starting from a leaf and walking to the root, we may handle point deactivations and point activations (Algorithm 6 and Algorithm 7). We say that node $v$ needs to handle a point insertion (or deletion) if for some child node $u$ of $v$ the set $L^{\prime}(u)$ undergoes a point insertion (or deletion). Both algorithms consist of procedures described in Algorithm 8 and Algorithm 9, in a bottom-up manner.
@@
First, assume, without loss of generality, that node $v\in T$ needs to handle a point insertion of $a\in A$ (see Algorithm 8). First we need to extend $L(v)$ by $a$. If $A(v)\geq B(v)$ then before the insertion $L^{\prime}(v)$ was either empty or consisted points of $A$. In this case we may simply extend $L^{\prime}(v)$ by $a$ as well and ask the parent of $v$ to handle the insertion of $a$.
@@ 13 $L^{\prime}(v)\leftarrow L^{\prime}(v)\cup\{a\}$
We can describe an analogous process for the deletion of point $a$ from node $v$ (see Algorithm 9). We first set $L(v)=L(v)\setminus\{a\}$. If $a\in L^{\prime}(v)$ then we can simply set $L^{\prime}(v)=L^{\prime}(v)\setminus\{a\}$ and ask the parent of $v$ to handle the deletion of $a$. Otherwise, if $A(v)>B(v)$ before the update we can select an arbitrary $a^{\prime}$ point of $L^{\prime}(v)$ and set $\mu(a^{\prime})=\mu(a)$, delete $a$ from $\mu$, set $L^{\prime}(v)=L^{\prime}(v)\setminus\{a^{\prime}\}$ and ask the parent of $v$ to handle the deletion of $a^{\prime}$. If $B(v)\geq A(v)$ before the deletion then let $b=\mu(a)$, we, delete the value of $a$ from $\mu$, set $L^{\prime}(v)=L^{\prime}(v)\cup b$ and ask the parent of $v$ to handle the insertion of $b$.
@@
Dynamic algorithms Algorithm 6 and Algorithm 7 maintain the solution to the Euclidean bipartite matching w.r.t. tree metric $\mathcal{T}$.
@@
We prove the above claim by showing that the invariant Lemma 5.7 is maintained after every update.
@@ 12 $L^{\prime}(v)\leftarrow L^{\prime}(v)\setminus\{a\}$
Corollary 5.5 now follows directly from Lemma 5.9 and Lemma 5.10 and Theorem 4.2.
@@ 5.3 Geometric Transport
In this section we describe our implementation of the fully dynamic geometric transport problem. The algorithm is mostly identical to that of Section 5.2 for Euclidean bipartite matching. However, in the case of geometric transport we allow for point pairs with large weights to be inserted and deleted from the input, hence the recourse of the actual assignment describing the optimal solution can be large after each update. Therefore, our algorithm only maintains the cost of the optimal solution.
@@
Similarly to our other applications, the algorithm can either obtain $O_{\epsilon}(\log n)$-approximation in $\tilde{O}(n^{\epsilon}+d)$ update time or $\tilde{O}(1)$-approximation in $\tilde{O}(d)$ update time. Our dynamic implementation will maintain the exact cost of an optimal solution with respect to the tree embedding described by Theorem 4.2. Hence, Corollary 5.11 is a direct implication of Lemma 4.4 and Lemma 5.12.
@@
First, notice that we can trivially turn an instance of geometric transport into an instance of Euclidean bipartite matching through creating $w(v)$ copies for all leaves $v$ in $T$. Therefore, it is enough to show that the cost of the matching $\mu$ produced by the static algorithm Algorithm 5 of Section 5.2 is exactly $\tilde{\mu}=\sum_{v\in T}||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$.
Now fix some arbitrary optimal matching $\mu^{*}$ on this Euclidean matching instance. Let $\Pi$ and $\Pi^{*}$ correspond to the multi-set of edges described by the union of paths on $T$ connecting the matched points of $\mu$ and $\mu^{*}$ respectively. Fix an arbitrary edge between $v$ and its parent $u$. In $\Pi^{*}$ this edge must appear at least $||A_{w}(v)|-|B_{w}(v)||$ as that many points of $P(v)$ cannot be matched to each other. Similarly, due to Lemma 5.7 this edge will appear at most $||A_{w}(v)|-|B_{w}(v)||$ times.
The paths in $\Pi$ crossing this edge contribute at most $||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$ to $c(\mu)$. Similarly, the paths in $\Pi^{*}$ crossing this edge due to crossing it contribute at least $||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$ to $c_{T}(\mu^{*})$ (as $\mu^{*}$ might not match $\min\{|A_{w}(v)|,|B_{w}(v)|\}$ pairs from $P(v)$). Summing over all $v\in T$, we get $c(\mu)\leq\tilde{\mu}\leq c(\mu^{*})$. As $c(\mu)\geq c(\mu^{*})$, we finally conclude that $c(\mu)=c(\mu^{*})=\sum_{v\in T}||A_{w}(v)|-|B_{w}(v)||\cdot 2^{m-\ell(v)}$. ∎
@@ 5.4 Euclidean Matching
In this subsection we describe our dynamic algorithm for dynamic Euclidean matching. The algorithm is similar to that of Section 5.2 for the bipartite version of the problem.
@@
We will maintain a tree embedding of the input points using Theorem 4.2. Let $T$ be the $2$-HST maintained by the algorithm, where leaves correspond the points of $P$. We will maintain an exact solution to the problem with respect to the distance metric defined by $T$. Then Corollary 5.14 is a direct implication of Lemma 4.4 and Lemma 5.15.
@@ 5 $L^{\prime}(v)\leftarrow\emptyset$
Observe that the matching $\mu$ computed by the static algorithm satisfies 5.16 by construction. Now we will show that this condition is sufficient to prove optimality.
@@
If matching $\mu$ satisfies 5.16 then $\mu$ is a minimum cost matching of $P$ with respect to the tree metric $\mathcal{\tau}$.
@@
Consider an arbitrary node $v\in T$ with parent $u$. If $|P(v)|$ is odd then $\mu^{*}$ must match a point of $P(v)$ with some point outside of $P(v)$. Hence $\Pi^{*}$ contains $(u,v)$. By assumption $\mu$ only matches one point of $P(v)$ outside of $P(v)$, hence $\Pi$ contains $(u,v)$ once. If $|P(v)|$ is even then, $\mu$ matched all points of $P(v)$ with each other, hence $(u,v)\notin\Pi$. This implies that $c_{T}(\mu)\leq c_{T}(\mu^{*})$. ∎
@@
We will now describe how to dynamically maintain a matching $\mu$ satisfying 5.16 under Type 1) and Type 2) updates in worst-case $\tilde{O}(1)$ update time. For sake of simplicity we will assume that $|P|$ is even at all times, that is updates come in pairs.
@@ 15    $\textbf{Delete}(u,p^{\prime\prime})$
The recursion follows a leaf to root path and takes $O(1)$ time to execute for a given note hence runs in $O(m)=\tilde{O}(1)$ time. Furthermore, it ensures that from $P(v)$ only the up to one point in $L^{\prime}(v)$ is not matched with points in $P(v)$. Hence, the algorithm satisfies 5.16 and by 5.17 it maintains a minimum cost matching of $P$ with respect to the tree metric $\mathcal{\tau}$.
Appendix A Local Queries for Hash Functions.
In Lemma 4.4, we state that the $(\Gamma,\Lambda)$-hash $\varphi$ with diameter bound $\tau$ can evaluate $\varphi(B(p,\tau/\Gamma))$ in $O(\operatorname{poly}(d)\cdot|\varphi(B(p,\tau/\Gamma))|$ time. However, this part of the statement did not appear in [FJL+25], and here we provide a proof sketch.